# A New Look at Fiducial Inference

### v1

@@ Abstract
Since the idea of fiducial inference was put forward by Fisher, researchers have been attempting to place it within a rigorous and well motivated framework. It is fair to say that a general definition has remained elusive. In this paper we start with a representation of Bayesian posterior distributions provided by Doob that relies on martingales. This is explicit in defining how a true parameter value should depend on a random sample and hence an approach to “inverse probability” (Fisher, 1930). Taking this as our cue, we introduce a definition of fiducial inference that extends existing ones due to Hannig.
@@ 1 Introduction
Our aim is not to reproduce another history; our starting point is the paper of Hannig (2009), also Hannig et al. (2016), which introduces a general version of fiducial inference. Hannig’s 2009 paper gives a principle under which a true parameter value can be obtained from a function, an independent random variable, and a summary statistic representing the data. However, a sequential version of Hannig’s construction, to be explained shortly, and based on the idea of sampling a complete data set, can be made to coincide with the Doob definition of a true parameter value conditioning on data, Doob (1949). It is straightforward to make this strategy ‘prior free’ using similar ideas presented in Fong et al. (2023), leading to a general, sequentially coherent, definition of a fiducial distribution.
@@ 1.1 The set up
whenever it exists as a cumulative distribution function. The constraints are obvious, $F_{n}(t\mid\theta)$ must be monotone in $\theta$ and be 00 and $1$ for the relevant extreme values of $\theta$. The corresponding fiducial density function is given by $\pi(\theta\mid t_{n})=-\frac{\partial F_{n}}{\partial\theta}\,(t_{n}\mid\theta).$ This seems, and is, quite restrictive.
@@
Our extension of Hannig’s notion of a fiducial distribution is based on Doob, and defined sequentially by
@@
where the $Z_{m}$ and $T_{m}$ are independent r.v. Assuming $T_{m}$ has a limit $T_{\infty}$, to exist a.s., it is the $T_{\infty}$ which is a sample from our fiducial distribution. $T_{\infty}=\lim_{m\to\infty}T_{m}$.
@@
The Doob Fiducial distribution is the distribution of the limit $T_{\infty}$ which is constructed from (3) and started at the observed $m=n$ and $T_{n}=t_{n}$. Here the $(Z_{m})$ are an independent sequence of random variables and the $H_{m}$ a known sequence of functions.
1.2 Illustration
@@
We produce the sequence $(T_{m})_{m>n}$, where $T_{m}=m^{-1}\sum_{i=1}^{m}X_{i}$ with $X_{m+1}=Z_{m}\,T_{m}/a$, and the $(Z_{m})$ are independent $\mbox{Ga}(a,1)$ variables. Hence, $T_{m+1}=T_{m}(m+Z_{m}/a)/(m+1)$, so completing the term from Definition 1, we have $H_{m}(t,z)=t(m+z/a)/(m+1)$. The rationale and theory behind this strategy will be gone through in the remainder of the paper.
Let us now describe the layout of the paper. In section 2 we present some further background to the fiducial distribution of Fisher and provide an interpretation of it which relies on the bootstrap of Efron. This demonstrates a lack of rigor with the idea and indeed highlights Fisher’s fiducial distribution as a heuristic approach. We also show how Doob’s interpretation of Bayes can be used to define a well motivated fiducial distribution, the key to which is a sequential development of connections between samples and parameters. Section 3 provides the theoretical results which establish conditions under which the Doob fiducial distribution exists. Section 4 contains some illustrations to make clear the novelty and outline of the new fiducial distribution. These examples include a normal, exponential model and a Weibull and uniform model. Finally, section 6 concludes with a brief discussion.
@@ 2 Background
Before we provide further explanation on the Doob fiducial approach, we show how it is possible to interpret the Fisher fiducial approach as a parametric bootstrap in the spirit of Efron; see Efron (2012). Following the original ideas of Fisher, the construction of the fiducial distribution can be shown to be based on the notion that if $T_{n}(\theta)$ is a statistic produced by taking $n$ i.i.d. samples from $f(\cdot\mid\theta)$, and the observed $t_{n}$ is arising in the same way with $\theta^{*}$, then the inequality $T_{n}(\theta)\leq t_{n}(\theta^{*})=t_{n}$ is suggestive of $\theta<\theta^{*}$. Of course this is not true due to stochastic variation. But as a guide, Fisher uses this to derive
${\rm P}(T>\theta\mid t_{n})={\rm P}(T_{n}(\theta)\leq t_{n})=F_{n}(t_{n}\mid\theta),$
being $T$ the random counterpart of $\theta^{*}$. This can now be converted into the fiducial distribution as in (1).
@@
Such a fiducial distribution is well defined provided that $t_{n}<n$ and in such case is Beta$(1+t_{n},n-t_{n})$. One can see this recalling the binomial expansion of the regularized incomplete beta function, given for instance by Abramowitz and Stegun (1964) in their formula 6.6.4 on page 263, namely:
@@
which holds for $10<p<$ 1, $m=1,\dotsc,n$, $n=1,2,\dotsc$
We can achieve this by taking repeated bootstrap samples $x_{1:n}(\theta)$ as i.i.d. Ber$(\theta)$ for each $\theta$. Set $T_{n}(\theta)$ to be the $\sum_{i=1}^{n}x_{i}(\theta)$. Then $T_{n}(\theta)>t_{n}$ is evidence of $\theta>\theta^{*}$. So repeating this $B$ times, an Efron style bootstrap estimator for the distribution of $T$ is
$\Pi^{(B)}(\theta\mid t_{n})=B^{-1}\sum_{b=1}^{B}1(T_{n}^{(b)}(\theta)>t_{n}),$
@@
so that the fiducial distribution is Gamma$(1+t_{n},n)$. The second identity in (4) is a consequence of a well known relationship between the Gamma and the Poisson distribution (Johnson et al., 2005, formula 4.96, page 197), namely ${\rm P}(X\leq x)={\rm P}(Y\geq m)$ if $X$ is Gamma$(m,\beta)$, being $m$ an integer, $\beta>0$ and $Y$ is Poisson$(\beta x)$ with $x>0$.
@@
Hence, we see that both Fisher’s original definition and the Hannig extension have some interpretation concerns. On the other hand, Doob (1949) has adequately explained how $\theta^{*}$ can be recovered from an infinite sample, which is where we start our version of the fiducial distribution. So Doob showed, for identifiable parameters, that it is possible to write $\theta^{*}=H(X_{1:\infty})$. This is based on the fact that
[…]
So $\theta^{*}$ is available from the infinite random sample.
@@
for some $\widetilde{H}_{m}$. An appropriate choice for the conditional distribution of $X_{m+1:\infty}$ given $T_{m}=t_{m}$ is such that for every $N>m$ the conditional density of $x_{m+1:N}$ given $T_{m}=t_{m}$ is
$\prod_{l=m+1}^{N}f(x_{l}\mid t_{l-1}).$
A sequential version of (5) is therefore given by $T_{m+1}=H_{m}(T_{m},Z_{m})$ for some sequence of functions $(H_{m})$ where the assumption is that $Z_{m}$ is independent of $T_{m}$; i.e. the $X_{1:m}$.
@@ 3 General Models
In this section we consider the convergence of the sequence (3) to a limit random variable $T_{\infty}$. In the following we introduce the function $g$ for which it is possible to sample $X_{m+1}=g(T_{m},Z_{m})$ from the density function $f(\cdot\mid T_{m})$. For example, for the gamma model used in the Introduction, $g(t,z)=tz/a$.
@@
    $\begin{split}\sum_{m=1}^{\infty}&{\rm E}(g(T_{m},Z_{m})\mid T_{m}),\\ \sum_{m=1}^{\infty}&{\rm E}(g(T_{m},Z_{m})^{2}\mid T_{m}),\end{split}$ (7)
@@
A general multivariate parameter form for this, with one dimensional data, has, for $j=1,\ldots,d$, $T_{m+1,j}=H_{m,j}(T_{m},Z_{m}),$ where $T_{m}=(T_{m,1},\ldots,T_{m,d})$ and $Z_{m}$ is a one dimensional innovation variable for which $X_{m+1}=g(T_{m},Z_{m})$. Then $T_{m+1,j}=\varphi_{j}^{-1}(\varphi_{j}(T_{m,j})+g(T_{m},Z_{m})).$
@@ 3.1 Proof of Theorem 1
$\begin{split}\sum_{m=1}^{\infty}&{\rm E}(Y_{m}\mid\mathscr{F}_{m-1}),\\ \sum_{m=1}^{\infty}&{\rm E}(Y_{m}^{2}\mid\mathscr{F}_{m-1}),\end{split}$ (10)
@@
We are now ready to prove Theorem 1. Combining (3) and (6) we have that $\varphi(T_{m+1})-\varphi(T_{m})=g(T_{m},z_{m})$ for every $m\geq n$. Therefore,
$\begin{split}\varphi(T_{m})&=\varphi(T_{n})+\sum_{\ell=n}^{m-1}\{\varphi(T_{\ell+1})-\varphi(T_{\ell})\}\\ &=\varphi(T_{n})+\sum_{\ell=n}^{m-1}g(T_{\ell},z_{\ell}),\end{split}$ (12)
@@
For every $m\geq 1$, let $\mathscr{F}_{m}$ be the sigma-field generated by the $T_{1},Z_{1},\dotsc,Z_{m}$. Applying inductively the recursion (8), one can see that $T_{m}$ is a function of $T_{1},Z_{1},\dotsc,Z_{m-1}$, say $T_{m}=h_{m}(T_{1},Z_{1},\dotsc,Z_{m-1})$ for some function $h_{m}$, and therefore $T_{m}$ and $Z_{m}$ are independent. Hence, setting $\phi_{m}(t)={\rm E}(g(t,Z_{m})),$ for every $t$, by Lemma 1 we have that
$\begin{split}{\rm E}(g(T_{m},Z_{m})&\mid\mathscr{F}_{m-1})\\ &={\rm E}(g(h_{m}(T_{1},Z_{1},\dotsc,Z_{m-1}),Z_{m})\mid T_{1},Z_{1},\dotsc,Z_{m-1})\\ &=\phi_{m}(T_{m}),\end{split}$ (13)
@@
${\rm E}(g(T_{m},Z_{m})^{2}\mid\mathscr{F}_{m-1})={\rm E}(g(T_{m},Z_{m})^{2}\mid T_{m}).$
In other words, the two random series (7) are both equal to the $\sum_{m=1}^{\infty}{\rm E}(g(T_{m},Z_{m})\mid\mathscr{F}_{m}),$ and $\sum_{m=1}^{\infty}{\rm E}(g(T_{m},Z_{m})^{2}\mid\mathscr{F}_{m})$, respectively. Since the two random series converge almost surely, we can apply Corollary 1 setting $Y_{m}=g(T_{m},Z_{m})$. This proves that the random series $\sum_{m=n}^{\infty}g(T_{m},Z_{m})$ converges almost surely. Therefore, by (12), $\varphi(T_{m})$ converges almost surely to some random variable $W$. Being $\varphi^{-1}$ continuous by assumption, we can apply the continuous mapping theorem to obtain that $T_{m}$ converges almost surely to $\varphi^{-1}(W)$ and the proof of Theorem 1 is complete.
@@ 4.1 Normal model
A normal model with unknown mean $\theta$ and known variance $\sigma$, and observed $y_{1:n}$, so let $t_{n}=\bar{y}$. Then, according to the sampling scheme for $Y_{n+1:\infty}$ where $Y_{m+1}\mid T_{m}\sim N(T_{m},\sigma^{2})$, it is that
$T_{m+1}=T_{m}+\frac{\sigma}{m+1}\,Z_{m},\quad m\geq n,$
where the $(Z_{m})$ are independent standard normal. So
@@
For a multivariate case, consider the case when $\theta$ and $\sigma$ are unknown. The start is $T_{n}$ to be the sample mean and $S_{n}^{2}$ to be the sample variance. Then for a general $m$, $Y_{m+1}=T_{m}+S_{m}\,Z_{m}$ where $(Z_{m})$ is a sequence of independent standard normal random variables. The updates are $T_{m+1}=T_{m}+S_{m}\,Z_{m}/(m+1)$ and
$S_{m+1}^{2}=S_{m}^{2}+S_{m}^{2}\left\{\frac{Z_{m}^{2}}{m+1}-\frac{1}{m}\right\}.$
So

### v2

@@ Abstract
Since the idea of fiducial inference was put forward by Fisher, researchers have been attempting to place it within a rigorous and well motivated framework. It is fair to say that a general definition has remained elusive. In this paper we start with a representation of Bayesian posterior distributions provided by Doob that relies on martingales. This is explicit in defining how a true parameter value should depend on a random sample and hence an approach to “inverse probability” as originally conceived by Fisher. Taking this as our cue, we introduce a definition of fiducial inference that can be regarded as general.
@@ 1 Introduction
Our aim is not to reproduce another history; our starting point is the paper of Hannig (2009), also Hannig et al. (2016), which introduces a general version of fiducial inference. Hannig’s 2009 paper gives a principle under which a true parameter value can be obtained from a function, an independent random variable, and a summary statistic representing the data. In this paper, we introduce a sequential version of Hannig’s construction, to be explained shortly, is based on the idea of sampling a complete data set, and which can be made to coincide with the Doob definition of a true parameter value conditioning on data; see Doob (1949). It is straightforward to make this strategy ‘prior free’ using the ideas presented in Fong et al. (2023), leading to a general, sequentially coherent, definition of a fiducial distribution.
In particular, we are able to consider the general framework used by Liang et al. (2025). This assumes observations $Y_{1:n}$, a vector of size $n$, is connected to covariates $X_{1:n}$, a $n\times p$ matrix, and parameter $\theta$, a $p$-dimensional vector, and independent random variable $Z_{1:n}$, via
$Y_{1:n}=H(X_{1:n},Z_{1:n},\theta),$ (1)
with inverse function $G$, such that $\theta=G(Y_{1:n},X_{1:n},Z_{1:n})$. See Assumption 1 and equation (1) in Liang et al. (2025). Our development is to recast (1) into a sequential one-step ahead form and to look directly at a sequence of statistics $(T_{m})_{m>n}$ for which we assign a distribution conditional on the observed statistic $T_{n}$. So (1) would become $T_{m}=H(X_{m},Z_{m},T_{m-1}),\quad m>n.$ In short, this is replacing the $\theta$ in (1) with estimator $T_{m-1}$. The limit of $T_{m}$ as $m\to\infty$ represents a posterior sample from our fiducial distribution, provided it exists, and it this existence which forms the theoretical part of the present paper. This approach mirrors the Doob (1949) definition of a posterior distribution using predictive distributions. The advantage here is that we do not require the need for an inverse function, which to the present day, has been one of the most important and yet severe restrictions on extending the applicability of fiducial inference.
@@ 1.1 The set up
whenever it exists as a cumulative distribution function. The constraints are obvious, $F_{n}(t\mid\theta)$ must be monotone in $\theta$ and be 0 and $1$ for the relevant extreme values of $\theta$. The corresponding fiducial density function is given by $\pi(\theta\mid t_{n})=-\partial F_{n}/\partial\theta\,(t_{n}\mid\theta).$ This seems, and is, quite restrictive.
@@
Our extension of Hannig’s notion of a fiducial distribution is based on Doob, and is defined sequentially by
@@
where the $Z_{m}$ and $T_{m}$ are independent r.v. Assuming $T_{m}$ has a limit $T_{\infty}$, to exist a.s., it is the $T_{\infty}$ which is a sample from our fiducial distribution; i.e. $T_{\infty}=\lim_{m\to\infty}T_{m}$. We can and do include covariate information later on, though for now we define the fiducial distribution without them:
@@
The Doob Fiducial distribution is the distribution of the limit $T_{\infty}$ which is constructed from (4) and started at the observed $m=n$ with $T_{n}=t_{n}$. Here the $(Z_{m})$ are an independent sequence of random variables and the $H_{m}$ a known sequence of functions.
1.2 Demonstrative illustration
@@
We produce the sequence $(T_{m})_{m>n}$, where $T_{m}=m^{-1}\sum_{i=1}^{m}X_{i}$ with $X_{m+1}=Z_{m}\,T_{m}/a$, and the $(Z_{m})$ are independent $\mbox{Ga}(a,1)$ variables. Hence, $T_{m+1}=T_{m}(m+Z_{m}/a)/(m+1)$, so completing the term from Definition 1, we have $H_{m}(t,z)=t(m+z/a)/(m+1)$. The fiducial distribution is the distribution of the limit $T_{\infty}$. The rationale and theory behind this strategy will be gone through in the remainder of the paper.
Let us now describe the layout of the paper. In section 2 we present some further background to the fiducial distribution of Fisher and provide an interpretation of it which relies on the bootstrap of Efron. This demonstrates a lack of rigor with the idea and indeed highlights Fisher’s fiducial distribution as a heuristic approach. We also show how Doob’s interpretation of Bayes can be used to define a well motivated fiducial distribution, the key to which is a sequential development of connections between samples and parameters. Section 3 provides the theoretical results which establish conditions under which the Doob fiducial distribution exists. Section 4 contains some illustrations to make clear the novelty and outline of the new fiducial distribution. These examples include a normal, exponential model and a Weibull and uniform model. Section 5 deals with nonparametric regression. An illustration related to logistic regression is provided. Finally, section 6 concludes with a brief discussion.
@@ 2 Background
Before we provide further explanation on the Doob fiducial approach, we show how it is possible to interpret the Fisher fiducial approach as a parametric bootstrap in the spirit of Efron; see Efron (2012). Following the original ideas of Fisher, the construction of the fiducial distribution can be shown to be based on the notion that if $T_{n}(\theta)$ is a statistic produced by taking $n$ i.i.d. samples from $f(\cdot\mid\theta)$, and the observed $t_{n}$ is arising in the same way with $\theta^{*}$, then the inequality $T_{n}(\theta)\leq t_{n}(\theta^{*})=t_{n}$ is suggestive of $\theta<\theta^{*}$. Of course this is not true due to stochastic variation. But as a guide, Fisher uses this to derive ${\rm P}(T>\theta\mid t_{n})={\rm P}(T_{n}(\theta)\leq t_{n})=F_{n}(t_{n}\mid\theta),$ being $T$ the random counterpart of $\theta^{*}$. This can now be converted into the fiducial distribution as in (2).
@@
[…]
@@
which holds for $0<p<1$, $m=1,\dotsc,n$, and $n=1,2,\dotsc$.
[…]
$\Pi^{(B)}(\theta\mid t_{n})=B^{-1}\sum_{b=1}^{B}1\bigg(T_{n}^{(b)}(\theta)>t_{n}\bigg),$
@@
[…]
@@
Hence, we see that both Fisher’s original definition and the Hannig extension have some interpretation concerns. On the other hand, Doob (1949) has adequately explained how $\theta^{*}$ can be recovered from an infinite sample, which is where we start our version of the fiducial distribution. So Doob showed, for identifiable parameters, that it is possible to write $\theta^{*}=H(X_{1:\infty})$. This is based on the fact that $\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}1(X_{i}\leq x)=F(x\mid\theta^{*})\,\mbox{a.s.}$ So $\theta^{*}$ is available from the infinite random sample.
@@
for some $\widetilde{H}_{m}$. An appropriate choice for the conditional distribution of $X_{m+1:\infty}$ given $T_{m}=t_{m}$ is such that for every $N>m$ the conditional density of $x_{m+1:N}$ given $T_{m}=t_{m}$ is $\prod_{l=m+1}^{N}f(x_{l}\mid t_{l-1}).$ A sequential version of (6) is therefore given by $T_{m+1}=H_{m}(T_{m},Z_{m})$ for some sequence of functions $(H_{m})$ where the assumption is that $Z_{m}$ is independent of $T_{m}$; i.e. the $X_{1:m}$.
@@ 3 General Models
In this section we consider the convergence of the sequence (4) to a limit random variable $T_{\infty}$. In the following we consider $H_{m}(t,z)=\varphi^{-1}(\varphi(t)+g_{m}(t,z))$. This is a very general expression for $H_{m}$. For example, for the gamma model $\varphi(t)=\log(t)$ and $g_{m}(t,z)=\log((m+z/a)/(m+1))$.
@@
    $\begin{split}\sum_{m=1}^{\infty}&{\rm E}(g_{m}(T_{m},Z_{m})\mid T_{m}),\\ \sum_{m=1}^{\infty}&{\rm E}(g_{m}(T_{m},Z_{m})^{2}\mid T_{m}),\end{split}$ (8)
@@
A general multivariate parameter form for this, with one dimensional data, has, for $j=1,\ldots,d$, $T_{m+1,j}=H_{m,j}(T_{m},Z_{m}),$ where $T_{m}=(T_{m,1},\ldots,T_{m,d})$ and $Z_{m}$ is a one dimensional innovation variable for which $T_{m+1,j}=\varphi_{j}^{-1}(\varphi_{j}(T_{m,j})+g_{m}(T_{m},Z_{m})).$
@@ 3.1 Proof of Theorem 1
$\sum_{m=1}^{\infty}{\rm E}(Y_{m}\mid\mathscr{F}_{m-1})\quad\mbox{and}\quad\sum_{m=1}^{\infty}{\rm E}(Y_{m}^{2}\mid\mathscr{F}_{m-1}),$ (11)
@@
We are now ready to prove Theorem 1. Combining (4) and (7) we have that $\varphi(T_{m+1})-\varphi(T_{m})=g_{m}(T_{m},z_{m})$ for every $m\geq n$. Therefore,
$\varphi(T_{m})=\varphi(T_{n})+\sum_{\ell=n}^{m-1}\{\varphi(T_{\ell+1})-\varphi(T_{\ell})\}=\varphi(T_{n})+\sum_{\ell=n}^{m-1}g_{\ell}(T_{\ell},z_{\ell}),$ (13)
@@
For every $m\geq 1$, let $\mathscr{F}_{m}$ be the sigma-field generated by the $T_{1},Z_{1},\dotsc,Z_{m}$. Applying inductively the recursion (9), one can see that $T_{m}$ is a function of $T_{1},Z_{1},\dotsc,Z_{m-1}$, say $T_{m}=h_{m}(T_{1},Z_{1},\dotsc,Z_{m-1})$ for some function $h_{m}$, and therefore $T_{m}$ and $Z_{m}$ are independent. Hence, setting $\phi_{m}(t)={\rm E}(g_{m}(t,Z_{m})),$ for every $t$, by Lemma 1 we have that
$\begin{split}{\rm E}(g_{m}(T_{m},Z_{m})&\mid\mathscr{F}_{m-1})\\ &={\rm E}(g_{m}(h_{m}(T_{1},Z_{1},\dotsc,Z_{m-1}),Z_{m})\mid T_{1},Z_{1},\dotsc,Z_{m-1})\\ &=\phi_{m}(T_{m}),\end{split}$ (14)
@@
${\rm E}(g_{m}(T_{m},Z_{m})^{2}\mid\mathscr{F}_{m-1})={\rm E}(g_{m}(T_{m},Z_{m})^{2}\mid T_{m}).$
So the two random series (8) are both equal to the $\sum_{m=1}^{\infty}{\rm E}(g_{m}(T_{m},Z_{m})\mid\mathscr{F}_{m}),$ and $\sum_{m=1}^{\infty}{\rm E}(g_{m}(T_{m},Z_{m})^{2}\mid\mathscr{F}_{m})$, respectively. Since the two random series converge almost surely, we can apply Corollary 1 setting $Y_{m}=g_{m}(T_{m},Z_{m})$. This proves that the random series $\sum_{m=n}^{\infty}g_{m}(T_{m},Z_{m})$ converges almost surely. Therefore, by (13), $\varphi(T_{m})$ converges almost surely to some random variable $W$. Being $\varphi^{-1}$ continuous by assumption, we can apply the continuous mapping theorem to obtain that $T_{m}$ converges almost surely to $\varphi^{-1}(W)$ and the proof of Theorem 1 is complete.
@@ 4.1 Normal model
A normal model with unknown mean $\theta$ and known variance $\sigma$, and observed $y_{1:n}$, so let $t_{n}=\bar{y}$. Then, according to the sampling scheme for $Y_{n+1:\infty}$ where $Y_{m+1}\mid T_{m}\sim N(T_{m},\sigma^{2})$, it is that $T_{m+1}=T_{m}+Z_{m}\,\sigma/(m+1)\,,\quad m\geq n,$ where the $(Z_{m})$ are independent standard normal. So
@@
For a multivariate case, consider the case when $\theta$ and $\sigma$ are unknown. The start is $T_{n}$ to be the sample mean and $S_{n}^{2}$ to be the sample variance. Then for a general $m$, $Y_{m+1}=T_{m}+S_{m}\,Z_{m}$ where $(Z_{m})$ is a sequence of independent standard normal random variables. The updates are $T_{m+1}=T_{m}+S_{m}\,Z_{m}/(m+1)$ and $S_{m+1}^{2}=S_{m}^{2}+S_{m}^{2}\left\{Z_{m}^{2}/(m+1)-1/m\right\}.$ So